OpenStudy (anonymous):

simplify (3x^4y^8/6xy^2)^-4

OpenStudy (saifoo.khan):

Flip.

OpenStudy (saifoo.khan):

\[\large (\frac{6xy^2}{3x^4y^8} )^ 4\]Simplify..

OpenStudy (anonymous):

sorry cant read the expression send again

OpenStudy (anonymous):

what is the answer simplified?

OpenStudy (saifoo.khan):

\[\Large \frac{16}{x^{12}y^{24}}\]

OpenStudy (saifoo.khan):

can you read it?

OpenStudy (anonymous):

how did yoyget 16 on top can you show me the steps

OpenStudy (saifoo.khan):

2^4 = 16

OpenStudy (saifoo.khan):

we got 2 after cancelling out 6/3.